OpenStudy (anonymous):

Consider the vector v=(5,8). What is the angle between v=(5,8) and the positive x-axis?

OpenStudy (anonymous):

you can consider x- axis is a vector u = (1,0) and find the angle between them by formula theta = ..... do you know that formula?

OpenStudy (anonymous):

Nope, all i know is the quesstion. I just learned this tsuff

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

\[\theta = \cos^-(\frac{ u.v}{ ||u||||v||})\]

OpenStudy (anonymous):

that' it

OpenStudy (anonymous):

You would bput the numbers in with the equation but which number goes in which spot?

OpenStudy (anonymous):

oh, those are not numbers, those are vectors, and from numerator, vector u dot vector v. in denominator, ||u|| is the length of vector u and ||v|| is the length of vector v. you have formulas to figure out all of them

OpenStudy (anonymous):

So ino order to do this you will have to find the distance of V withc would be 5^2+*^2 right?

OpenStudy (anonymous):

\[||v||= \sqrt{5^2 +8^2}\]. you forgot sqrt

OpenStudy (anonymous):

length will be \[\sqrt{89}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then length of u =1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so how would we distribute for equation?

OpenStudy (anonymous):

how about u dot v at numerator?

OpenStudy (anonymous):

\[\frac{ \sqrt{89}\times1 }{ \left| 1 \right| }\left| \sqrt{89} \right|\]

OpenStudy (anonymous):

Not sure on bottom

OpenStudy (anonymous):

nope, u dot v = < (1,0), (5,8)> = 1*5 + 0*8 = 5 . that's just numerator, your denominator is sqr (89) so combine them you get \[\theta = \cos^- \frac{ 5 }{ \sqrt{89} }\]

OpenStudy (anonymous):

you have to use calculator to find the theta if you want answer in number, if not, that 's your answer, it's ok, too.

OpenStudy (anonymous):

For the exact number I got 57.99461679

OpenStudy (anonymous):

I don't do that part, trust yourself from that part.

OpenStudy (anonymous):

Alright, Thank you for your help!

OpenStudy (anonymous):

np